OpenStudy (anonymous):

The volume of a square based pyramid is 81 cubic meters. What is the volume of the cube with the same side length and height?

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

\[V_{\text{pyramid}}=\frac{bh}{3}\] \[b=lw\] \[\text{for a cube}\qquad l=w=h=s\] \[V_{\text{square pyramid}}=\frac{s^3}{3}\] \[V_{\text{cube}}={s^3}\]

OpenStudy (anonymous):

what ?

OpenStudy (unklerhaukus):